rationalize denominator ___5__
sq rt 3
___5___
2+sq rt 3
I am not getting this...I have like five more...and I just need to see them worked out so I can finish this part. Much Mahalo.

Answer by edjones(8007)

We are trying to get an integer in the denominator so we will have a rational number, that is all. What can you multiply times the denominator to give an integer? 2-sqrt(3) in this case. (remember: (a-b)(a+b)=a^2-b^2
5/(2+sqrt(3)) * (2-sqrt(3))/(2-sqrt(3))
=5(2-sqrt(3))/(4-3)
=10-5sqrt(3)
.
Ed